OpenStudy (anonymous):

Can someone please help me?!?! Simplify 4g+3(g+6)=

OpenStudy (anonymous):

first use the distributive law to remove the parentheses and write \[4g+3g+3\times 6=4g+3b+18\] then combine like terms,

OpenStudy (anonymous):

\(4g+3g=7g\) just like 4 dollars and 3 dollars is 7 dollars so your answer is \[7g+18\] clear more or less?